OpenStudy (anonymous):

what is the denominator of the simplified form of (1/2)-(3/4x-2) a.)2 b.)x-2 c.)2x-1

OpenStudy (campbell_st):

get the common denominator \[\frac{1}{2} -\frac{1}{2(2x -1)} = \frac{2x -1 -1}{2(2x-1} = \frac{2(x -1)}{2(2x -1)}\] cancel the common factor of 2 then you will have the simplified form all you need to do is identify the denominator